the length of a rectangle is 2 meters less than 5 times the width. if the perimeter of the rectangle is 92 meters, what is the area of the rectangle?
Answer by mananth(16946)

the length of a rectangle is 2 meters less than 5 times the width. if the perimeter of the rectangle is 92 meters, what is the area of the rectangle?
...
width =x
length = 5x-2
2(x+5x-2)=92
2(6x-2)=92
/2
6x-2=92/2
6x-2=46
+2
6x=48
/6
x=8m the width
length = 5x-2
40-2
=38 m
..
m.ananth@hotmail.ca